Week of 160222

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Alexei (LHCb), Maarten (ALICE), Maria A (WLCG), Julia (WLCG), Jose (DB), Andrea (FTS)
  • remote: David (ATLAS), Christoph (CMS), Francesco (CNAF), Rolf (IN2P3), Sang Un (KISTI), Christian (NDGF), Onno (NL-T1), Kyle (OSG), Pepe (PIC), Tiju (RAL), Di Qing (TRIUMF)

Experiments round table:

  • ATLAS reports (raw view) -
    • Reprocessing expected to finish in the next day or two
    • Taiwan cvmfs stratum 1 looks fixed
    • A campaign of cleaning old data from CERN and T1 tapes will start soon, sites will be contacted by ATLAS

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Data Processing
      • Stripping 24 is almost at 100%, the merging productions are following slowly. Validation of stripping for Turbo and TurCal
      • MC and User jobs

Sites / Services round table:

  • ASGC:
  • BNL:
  • CNAF: NTR
  • FNAL:
  • GridPP:
  • IN2P3: NTR
  • JINR:
  • KISTI: Downtime from midnight CEST to update glibc
  • KIT:
  • NDGF: Problem with a tape robot, some data might be unavailable. Upgrade on Thursday to fix it.
  • NL-T1: Upgrade of dCache and glibc
  • NRC-KI:
  • OSG: Several tickets for CMS T2 resources, handling them from this morning.
  • PIC: Back to full computing capacity after the fix of the critical vulnerability in glibc performed on Friday and during the weekend.
  • RAL: Patch glibc today for grid services and batch, tomorrow for FTS.
  • TRIUMF: Plan to do tape library upgrade from 17:00 February 24 to 08:00 on February 25 (UTC time) at TRIUMF. During the upgrade, reading from tape system is not available, however writing to tape still works, and disk system is not affected too.

  • CERN batch and grid services:
    • FTS Service upgrade to v 3.4.2 on Wed 24 from 9:30 to 10:30 CET. OTG:0028791 and on GOCDB. During the upgrade period nodes will be rebooted as well for the new glibc version so clients should expects some interruptions.
  • Storage: NTR
  • Databases: NTR
  • GGUS:
  • Grid Monitoring:
  • MW Officer: NTR

AOB:

  • From the 29th Feb the WLCG Operation meeting will be held only on Monday (the Thursday meeting will be cancelled). In addition sites are asked to pre-fill their report or to send an email to the SCOD list.

Thursday

Attendance:

  • local: Luca (SCOD+Storage), Maria A, Alexey (LHCb), Maarten (ALICE), Jiri (ATLAS), Laurence (Batch), Andrei (DB)
  • remote: Michael (BNL), Matteo (CNAF), David (FNAL), Rolf (IN2P3), Christian (NDGF), Andrew (Nikhef), Kyle (OSG), Tiju (RAL)

Experiments round table:

  • ATLAS reports (raw view) -
    • Amazon 100k test with BNL - starting up more jobs today after solving configuration issues
    • VOMS membership and AUP signature expiration OTG:0028871
      • Followed up by the CERN Batch Team

  • CMS reports (raw view) -
    • No major issues
    • Running at capacity
    • Started deployment of multicore at T2's, first set of 14 sites was asked to provide multicore resources, tracked in GGUS, will be focus of tomorrow's CMS Computing meeting (sites welcome, as usual).

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Data Processing
      • Active pre-staging
      • MC and User jobs

Sites / Services round table:

  • ASGC:
  • BNL:
    • NTR
    • Update on ATLAS 100k test: the setup of the computing capacity is complete, waiting now for the full commissioning of the storage element inside AGIS
  • CNAF: NTR
  • FNAL: Downtime yesterday to fix glibc vulnerability
  • GridPP:
  • IN2P3: NTR
  • JINR:
  • KISTI:
  • KIT:
  • NDGF: Downtime today to fix the robot issue. Tomorrow there will be 1h downtime to reboot a DB server and to update dCache
  • NL-T1: NTR
  • NRC-KI:
  • OSG: NTR
  • PIC:
  • RAL: NTR
  • TRIUMF:

  • CERN batch and grid services: NTR
  • CERN storage services: NTR
  • Databases:
    • ATLAS ATLR database patching and migration to new hardware, OTG:0028797
  • GGUS:
  • Grid Monitoring:
  • MW Officer:

AOB:

  • The CERN Batch team was asked to report in addition to the batch resources also for the OpenStack infrastructure
  • This was the last Thursday WLCG operation meeting, from next week this meeting will be held only on Monday.
    In addition sites are asked to pre-fill their report or to send an email to the SCOD list.

Status of deployment to fix last SV:

Site Status
ASGC DONE
BNL DONE
CNAF DONE
FNAL DONE
IN2P3 DONE
JINR ?
KISTI DONE
KIT DONE
NDGF DONE
NL-T1 DONE
NRC-KI ?
PIC DONE
RAL DONE
TRIUMF DONE
Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2016-02-25 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback